OpenStudy (anonymous):

Solve using the addition and multiplication principles. Be sure to show all of your work. -10 - 4b > 6 Solve using the addition and multiplication principles. Be sure to show all of your work. -10 - 4b > 6 @Mathematics

OpenStudy (anonymous):

give it a shot first.....

OpenStudy (anonymous):

my answer is 64b -10-4b>6 -10-6= -16 -16*4= 64b

OpenStudy (anonymous):

you have signs wrong, first step add 10 to both sides 10 + 6 = 16 not -16

OpenStudy (anonymous):

i was right though it was jux positive not negative

OpenStudy (anonymous):

-10 - 4b > 6 -4b > 16 b > -4

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

then i subtract -4b>16 = -4

OpenStudy (anonymous):

im confused bout the last part